If the total revenue of 7 units of a commodity is Birr 40 and the average

revenue of 8 units is Birr 5, what is the marginal revenue of the 8th

unit?


1
Expert's answer
2021-06-02T12:30:55-0400

Solution:

Total Revenue = P "\times" Q

Total revenue of 7 units (TR1) = 40

 

Average revenue = "\frac{TR}{ Q}"

Average revenue of 8 units = 5


5 = "\frac{TR}{ 8}"

 

Total revenue of 8 units (TR2) = 8 "\times" 5 = 40

 

Marginal revenue = "\frac{\triangle TR}{\triangle Q}"

MRn = TRn – TRn-1

Marginal revenue = TR1 – TR2

Marginal revenue = 40 – 40 = 0

Marginal revenue = 0